OpenStudy (anonymous):

The point P(6, 1) lies on the curve y = √x - 5. If Q is the point (x, √x - 5), use a scientific calculator to find the slope of the secant line PQ (correct to six decimal places) for the following values of x. x=5.5

OpenStudy (anonymous):

The points are \[ P=(6,1)\\ Q=(5.5, \sqrt {.5}) \]

OpenStudy (anonymous):

Slope \[ \frac { 1 - \sqrt{.5}}{6-5.5}=0.585786 \]